### v1

@@ 1 Introduction
All-optical neural networks (AONNs) represent a cutting-edge computational architecture that employs optical signals for end-to-end data processing, utilizing coherent light diffraction to perform spatial light computing and inference[1, 2, 3, 4, 5]. In contrast to conventional electronic neural networks, AONNs achieve orders-of-magnitude faster processing speeds by exploiting the inherent advantages of optical signals[6, 7, 8, 9], including low transmission loss and high bandwidth. These characteristics make them well-suited for high-throughput, low-latency inference tasks[10, 11, 12]. In electronic neural networks, nonlinear activation [13, 14, 15, 16]functions play a crucial role by enabling the model to learn abstract and discriminative feature representations through nonlinear transformations[17, 18]. This nonlinearity enhances the network’s ability to capture complex patterns in the input data and significantly improves feature separability, leading to higher accuracy in classification tasks [19, 20].
@@
However, the development of practical AONNs faces a critical challenge: the implementation of efficient all-optical nonlinear activation functions that can match the performance of their electronic counterparts while maintaining the energy efficiency and parallelism inherent to optical systems[21, 22]. Current approaches struggle to achieve both low power consumption and high nonlinearity simultaneously, often requiring either bulky atomic systems [23, 24] or high-intensity optical fields that introduce excessive energy dissipation. Moreover, most existing solutions lack the programmability needed to support diverse activation functions (e.g., ReLU, sigmoid) required for modern neural network architectures, and their integration with scalable photonic platforms remains a significant hurdle for large-scale deployment [25, 26, 27, 28, 29] . These limitations collectively constrain AONNs from realizing their full potential in complex machine learning tasks.
@@
The fundamental challenge in advancing AONNs lies in enabling multi-layer optical coupling and increasing the network’s tunable degrees[30, 31, 32] of freedom (or neuron count) without depending on conventional nonlinear optical components. Thus, recent research has explored several strategies to realize nonlinear computation through linear optical processes[33, 34]. One such approach employs multi-scattering cavities[35], where optical scattering generates input-dependent speckle patterns that passively mimic nonlinear transformations. However, this method suffers from uncontrollable noise and significant optical attenuation, necessitating high-power laser excitation. Another strategy involves dynamic diffractive layers[36], where iterative phase modulation enables superposition-based nonlinear approximations. While this reduces laser power requirements, it introduces hardware limitations due to the need for continuous parameter updates, leading to potential dynamic errors and speed constraints. Although both methods improve neuron density and computational accuracy, their respective trade-offs in controllability and system stability remain critical challenges for real-world implementation.
@@ 2 SP-AONNs framework
In the following, we demonstrate the advantages of parallel architecture by analyzing its neuron capacity enhancement. In optical neural networks, the number of neurons is determined by the independently tunable degrees of freedom within the network[36]. The input neuron count is defined by the image resolution (e.g., $1000\times 1000$ pixels), while the output neuron count depends on the output dimensions (e.g., $150\times 150$ pixels). For the SP-AONNs architecture with N parallel layers, the neuron count in the phase modulation layer comprises both intra-network components, representing independent degrees of freedom within each parallel sub-network that scale linearly with N, and inter-network coupling terms arising from coherent interference between sub-networks that scale quadratically with N (orange curve in Fig. 1d),
@@
A central advantage of the SP-AONNs framework in comparison to its series counterpart, lies in the fact that the SP-AONNS features an inherently coherence-induced nonlinear computing mechanism. To appreciate this point, let us consider the scalar diffraction theory[37, 38, 39], which tells the output light field $A_{i}$ of $i$-th sub-network is expressed as:
@@ 4 Discussion
Furthermore, the SP-AONNs framework exhibits superior noise resistance compared to serial architectures due to independent sub-network operation. We evaluated a trained 4-layer sub-network model by introducing phase noise ($0~-~\delta\cdot 2\pi$, $\delta\in[0,1]$) to each phase modulation layer, measuring classification accuracy across datasets (Fig. 4). For Digit MNIST, serial networks degraded to 10% accuracy (complete loss of classification capability) at $0.8\pi$ phase noise, whereas parallel networks maintained 83% accuracy under identical noise conditions. As shown in Fig.4, classification accuracy remains stable until phase noise intensity exceeds $\pi$. Since phase noise is a common problem in real-world optical neural networks, the SP-AONNs framework provides an effective way to maintain accuracy. In addition, incoherent stacking is often encountered in practical implementations of optical neural networks, since real-world light sources may be incoherent and detectors inherently measure optical intensity rather than field amplitude. Under such conditions, the spatially parallel architecture remains effective, although its accuracy is lower than that of coherent parallel architectures due to the absence of coherence-induced nonlinearity. Nevertheless, it still surpasses conventional serial architectures, highlighting the intrinsic advantage of the parallel framework (see supplementary material for details).
@@ 5 Conclusion
Furthermore, the parallel architecture’s non-accumulative error characteristic is a critical advantage, as it can effectively mitigate noise propagation throughout the optical system, thereby significantly improving overall noise tolerance. Future research directions will explore the integration of partially coherent light sources and other coherence-management techniques to further improve this inherent robustness [40, 41, 42]. For partially coherent light, the coherent components from different parallel channels can still interfere, preserving the nonlinear activation function’s effect. By integrating nonlinear activation functions with partially coherent light sources, we anticipate enhanced network robustness and improved target recognition accuracy. These advancements are essential for transitioning laboratory-scale demonstrations into practical, real-world applications where environmental stability is unpredictable. Thus, this work paves the way for the development of more reliable, efficient, and deployable optical computing systems for next-generation artificial intelligence.

### v2

@@ 1 Introduction
All-optical neural networks (AONNs) represent a cutting-edge computational architecture that employs optical signals for end-to-end data processing, utilizing coherent light diffraction to perform spatial light computing and inference[lin2018all, kues2017chip, liu2022programmable, zhou2021large, kulce2021all]. In contrast to conventional electronic neural networks, AONNs achieve orders-of-magnitude faster processing speeds by exploiting the inherent advantages of optical signals[nips1, nips2, nips3, nips4], including low transmission loss and high bandwidth. These characteristics make them well-suited for high-throughput, low-latency inference tasks[shastri2021photonics, wetzstein2020inference, feldmann2019all]. In electronic neural networks, nonlinear activation [nonline1, nonline2, nonline3, nonline4]functions play a crucial role by enabling the model to learn abstract and discriminative feature representations through nonlinear transformations[lecun2015deep, wang2023image]. This nonlinearity enhances the network’s ability to capture complex patterns in the input data and significantly improves feature separability, leading to higher accuracy in classification tasks [krizhevsky2012imagenet, goodfellow2016deep].
@@
However, the development of practical AONNs faces a critical challenge: the implementation of efficient all-optical nonlinear activation functions that can match the performance of their electronic counterparts while maintaining the energy efficiency and parallelism inherent to optical systems[hughes2018training, shen2017deep]. Current approaches struggle to achieve both low power consumption and high nonlinearity simultaneously, often requiring either bulky atomic systems [zuo2019all, zuo2021scalability] or high-intensity optical fields that introduce excessive energy dissipation. Moreover, most existing solutions lack the programmability needed to support diverse activation functions (e.g., ReLU, sigmoid) required for modern neural network architectures, and their integration with scalable photonic platforms remains a significant hurdle for large-scale deployment [multi1, multi2, multi3, multi4, multi5] . These limitations collectively constrain AONNs from realizing their full potential in complex machine learning tasks.
@@
The fundamental challenge in advancing AONNs lies in enabling multi-layer optical coupling and increasing the network’s tunable degrees[optnon1, optnon2, optnon3] of freedom (or neuron count) without depending on conventional nonlinear optical components. Thus, recent research has explored several strategies to realize nonlinear computation through linear optical processes[optline2, opyline1]. One such approach employs multi-scattering cavities[xia2024nonlinear], where optical scattering generates input-dependent speckle patterns that passively mimic nonlinear transformations. However, this method suffers from uncontrollable noise and significant optical attenuation, necessitating high-power laser excitation. Another strategy involves dynamic diffractive layers[yildirim2024nonlinear], where iterative phase modulation enables superposition-based nonlinear approximations. While this reduces laser power requirements, it introduces hardware limitations due to the need for continuous parameter updates, leading to potential dynamic errors and speed constraints. Although both methods improve neuron density and computational accuracy, their respective trade-offs in controllability and system stability remain critical challenges for real-world implementation.
@@ 2 SP-AONNs framework
In the following, we demonstrate the advantages of parallel architecture by analyzing its neuron capacity enhancement. In optical neural networks, the number of neurons is determined by the independently tunable degrees of freedom within the network[yildirim2024nonlinear]. The input neuron count is defined by the image resolution (e.g., $1000\times 1000$ pixels), while the output neuron count depends on the output dimensions (e.g., $150\times 150$ pixels). For the SP-AONNs architecture with N parallel layers, the neuron count in the phase modulation layer comprises both intra-network components, representing independent degrees of freedom within each parallel sub-network that scale linearly with N, and inter-network coupling terms arising from coherent interference between sub-networks that scale quadratically with N (orange curve in Fig. 1d),
@@
A central advantage of the SP-AONNs framework in comparison to its series counterpart, lies in the fact that the SP-AONNS features an inherently coherence-induced nonlinear computing mechanism. To appreciate this point, let us consider the scalar diffraction theory[theo1, theo2, theo3], which tells the output light field $A_{i}$ of $i$-th sub-network is expressed as:
@@ 4 Discussion
Furthermore, the SP-AONNs framework exhibits superior noise resistance compared to serial architectures due to independent sub-network operation. We evaluated a trained 4-layer sub-network model by introducing phase noise ($0~-~\delta\cdot 2\pi$, $\delta\in[0,1]$) to each phase modulation layer, measuring classification accuracy across datasets (Fig. 4). For Digit MNIST, serial networks degraded to 9.95% accuracy (complete loss of classification capability) at $1.0\pi$ phase noise, whereas parallel networks maintained 94.07% accuracy under identical noise conditions. As shown in Fig.4, classification accuracy remains stable until phase noise intensity exceeds $2\pi$. Since phase noise is a common problem in real-world optical neural networks, the SP-AONNs framework provides an effective way to maintain accuracy.
@@ 5 Conclusion
Furthermore, the parallel architecture’s non-accumulative error characteristic is a critical advantage, as it can effectively mitigate noise propagation throughout the optical system, thereby significantly improving overall noise tolerance. Future research directions will explore the integration of partially coherent light sources and other coherence-management techniques to further improve this inherent robustness [qin2025all, dong2024partial, jia2024partially]. For partially coherent light, the coherent components from different parallel channels can still interfere, preserving the nonlinear activation function’s effect. By integrating nonlinear activation functions with partially coherent light sources, we anticipate enhanced network robustness and improved target recognition accuracy. These advancements are essential for transitioning laboratory-scale demonstrations into practical, real-world applications where environmental stability is unpredictable. Thus, this work paves the way for the development of more reliable, efficient, and deployable optical computing systems for next-generation artificial intelligence.